# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 7 Conclusion
  • [26] Wei Shi, Qing Ling, Kun Yuan, Gang Wu, and Wotao Yin. On the linear convergence of the admm in decentralized consensus optimization. Signal Processing, IEEE Transactions on, 62, 07 2013.
@@
  • [29] A. B. Taylor, J. M. Hendrickx, and F. Glineur. Exact worst-case performance of first-order methods for composite convex optimization. SIAM Journal on Optimization, 27, 12 2015.
  • [30] A. B. Taylor, J. M. Hendrickx, and F. Glineur. Smooth strongly convex interpolation and exact worst-case performance of first-order methods. Mathematical Programming, 161, 02 2015.
@@ A.1 When all agents are equivalent
Theorem 7 characterizes the settings in which we can formulate the performance computation of a distributed optimization algorithm as an SDP PEP which is totally independent of the number of agents $n$. In particular, we need all the agents to be equivalent (Assumption 1), so we can work with a fully symmetric solution $(F^{s},G^{s})$ (Corollary 6.1), composed of small repeating blocks $f_{A}$, $G_{A}$, and $G_{B}$. The resulting agent-dependent PEP, expressed with blocks $f_{A}$, $G_{A}$, and $G_{B}$ has a size independent of the number of agents $n$, but the value of $n$ can still appear in the PEP and impact its solution, as any other parameter of the problem. Using suited changes of variables, we can obtain a PEP formulation that is fully independent of $n$, under the condition stated in part 2 of Theorem 7. In this appendix, we show how we can explicitly build this agent-independent PEP formulation for many usual settings. For that purpose, we pass through the main components of a PEP, that we have introduced in Section 2, and detail how they can be written with the blocks $f_{A}$, $G_{A}$, and $G_{B}$ and which changes of variable get rid of $n$.
@@
Let $\mathcal{F}$ be a set of functions for which there exist linearly Gram-representable interpolation constraints. In a PEP for a distributed optimization method, with fully symmetrized solutions $(F^{s},G^{s})$ (77)-(81), the constraints $f_{i}\in\mathcal{F}$, for all $i\in\mathcal{V}$ can be expressed using a set of interpolation constraints that are agent-independently Gram-representable and that only involves $f_{A}$ and $G_{A}$.
@@
Each constraint $f_{i}\in\mathcal{F}$ can be written with a set of Gram-representable interpolation constraints involving function values of agent $i$ and scalar products between quantities (i.e. points and gradients) related to agent $i$. These are thus single-agent constraints that are all identical when we restrict to agent-symmetric PEP solution $(F^{s},G^{s})$ (77)-(81). Therefore, they can all be expressed with $f_{A}$ and $G_{A}$ using (95) and (96). ∎
@@
Concerning the consensus steps, Theorem 7 applies to the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, defined in Section 2. Necessary interpolation constraints for this set of matrices are given in (50), (51) and (52), from Theorem 4 applied to the consensus subspace $\mathcal{C}$. Proposition 13 below shows that these interpolation constraints are agent-independently Gram-representable, by expressing them in terms of $G_{C}$ and $G_{D}$. In principle, the compact PEP formulation could apply to a general set of averaging matrices $\mathcal{M}$ for which there exist Gram-representable interpolation constraints. Future work may include the description of other classes of averaging matrices, in particular, the extension to non-symmetric stochastic matrices with a bound on their singular value.
@@
The averaging matrix interpolation constraints (50), (51) and (52), from Theorem 4 applied to the consensus subspace $\mathcal{C}$, are agent-independently Gram-representable and can be expressed using $G_{C}$ and $G_{D}\in\mathbb{R}^{p\times p}$ as:
@@
Firstly, we express (50) as (132). Equation (50), when applied to the consensus subspace, means that the agent average is preserved, i.e. $\overline{X}=\overline{Y}$, which can be written with scalar products as:
@@
[…]
@@
$\frac{1}{n}\sum_{i=1}^{n}(U_{i}-\overline{U})^{T}(V_{i}-\overline{V})=e_{m}^{T}G_{D}e_{V}=(e_{Y}-\lambda^{-}e_{X})^{T}G_{D}(e_{Y}-\lambda^{+}e_{X})\preceq 0$
@@
Points common to all agents ($x^{*}$, $\overline{x}$, etc)
@@
$F=\begin{bmatrix}f_{1}^{T}&\dots&f_{n}^{T}\end{bmatrix},\quad\text{where $f_{i}=\quantity[f_{i}^{k}]_{k\in I}\in\mathbb{R}^{q}$ is a vector with the $q$ functions values of agent $i$,}$ (139)
$P=\begin{bmatrix}P_{1}&\dots&P_{n}\end{bmatrix}\quad\text{where $P_{i}\in\mathbb{R}^{d\times p}$ contains the $p$ vector variables related to agent $i$,}$ (140)
e.g. $P_{i}=\quantity[y_{i}^{k}\leavevmode\nobreak\ x_{i}^{k}\leavevmode\nobreak\ g_{i}^{k}]_{k\in I}$, and the Gram matrix $G\in\mathbb{R}^{np\times np}$ is thus defined as
@@
$\overline{x}_{i}=\overline{x}=\frac{1}{n}\sum_{j=1}^{n}x_{j}\quad\text{for all$i\in\mathcal{V}$},\qquad\text{ or }\qquad\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(x_{i}^{*})=0,\quad\text{with $x^{*}_{i}=x^{*}_{j}=x^{*}$ for all $i,j\in\mathcal{V}$},$ (142)
@@
Let $x^{*}$ be an optimal point for the decentralized optimization problem (1). The definition constraints for the common variable $x^{*}$ in a PEP restricted to fully symmetric solutions $(F^{s},G^{s})$ (77)-(81), given by the system-level stationarity constraint
@@
We consider a PEP for distributed optimization, restricted to fully symmetric solutions $(F^{s},G^{s})$ (77)-(81). The definition constraints for $\overline{x}$
@@
$\frac{1}{n}\sum_{i=1}^{n}\quantity(f_{i}(\overline{x}^{0})-f_{i}(x^{*}))$ $\leq R$ $\quantity(e_{f(\overline{x}^{0})}-e_{f(x^{*})})^{T}f_{A}\quantity(e_{f(\overline{x}^{0})}-e_{f(x^{*})})$ $\leq R$ (155)
@@
To obtain the two first reformulations, we can use relation (93) applied respectively to vectors $x_{i}^{0}-x^{*}$ and $\nabla f_{i}(x_{i}^{0})$
@@ A.2 With multiple equivalence classes of agents
Let $\mathcal{T}$ be a partition of the agent set $\mathcal{V}$ into $m$ equivalence classes of agents (see Definition 9). Considering the PEP restricted to symmetrized agent-class solutions $G^{s}_{\mathcal{T}}$ and $F^{s}_{\mathcal{T}}$ (105)-(110), we now detail how we can express all the PEP elements in terms of $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,m$), to be able to use this blocks as variable of the compact SDP PEP formulation. For the SDP condition $G^{s}\succeq 0$, we can be express it with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ using Lemma 9.1. For the other PEP elements, we need to adapt the reformulation techniques (92), (93), (94), (95), (96) and (129) to the situation where the agent set $\mathcal{V}$ is partitioned into $m$ subsets of equivalent agents $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{m}\}$. The terms that we can be expressed in terms of $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ are the following (for any variables $x_{i}$ and $y_{j}$ related to agents $i$ and $j$, including the case $x_{i}=y_{j}$):
@@
    where the first equality holds because $f_{i}(x_{i})=f_{j}(x_{j})$ for any $i,j\in\mathcal{V}_{u}$, by definition of the symmetric agent-class solution $F^{s}_{\mathcal{T}}$ (105). This implies that
    $\frac{1}{n}\sum_{i=1}^{n}f_{i}(x_{i})=\frac{1}{n}\sum_{u=1}^{m}\sum_{j\in V_{u}}^{n}f_{j}(x_{j})=\frac{1}{n}\sum_{u=1}^{m}n_{u}e_{f(x)}^{T}f_{A}^{u}=e_{f(x)}^{T}f_{A,\mathcal{T}},$ (159)
@@
    $\frac{1}{n}\sum_{i=1}^{n}x_{i}^{T}y_{i}=\frac{1}{n}\sum_{u=1}^{m}\sum_{j\in V_{u}}^{n}x_{j}^{T}y_{j}=\frac{1}{n}\sum_{u=1}^{m}n_{u}e_{x}^{T}G_{A}^{u}e_{y}=e_{x}^{T}(G_{A,\mathcal{T}})e_{y},$ (162)
@@
    $G_{A,\mathcal{T}}=\frac{1}{n}\sum_{u=1}^{m}n_{u}G_{A}^{u}.$ (163)
  • The average over the $n^{2}$ pairs of agents of the scalar products between two variables each related to any agent:
    $\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}x_{i}^{T}y_{j}$ $=\frac{1}{n^{2}}\sum_{u=1}^{m}\sum_{v=1}^{m}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{v}}x_{i}^{T}y_{j}=\frac{1}{n^{2}}\sum_{u=1}^{m}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{u}}x_{i}^{T}y_{j}+\frac{1}{n^{2}}\sum_{u=1}^{m}\sum_{v\neq u}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{v}}x_{i}^{T}y_{j},$ (164)
    $\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}x_{i}^{T}y_{j}$ $=e_{x}^{T}\quantity(\frac{1}{n^{2}}\sum_{u=1}^{m}\bigl{(}n_{u}G_{A}^{u}+n_{u}(n_{u}-1)G_{B}^{u}\bigr{)}+\frac{1}{n^{2}}\sum_{u=1}^{m}\sum_{v\neq u}n_{u}n_{v}\,G_{E}^{uv})e_{y}=e_{x}^{T}(G_{C,\mathcal{T}})e_{y},$ (165)
@@
    $G_{C,\mathcal{T}}=\frac{1}{n^{2}}\quantity(\sum_{u=1}^{m}n_{u}(G_{A}^{u}+(n_{u}-1)G_{B}^{u})+\sum_{u=1}^{m}\sum_{v\neq u}n_{u}n_{v}\,G_{E}^{uv}).$ (166)
@@
    $\frac{1}{n}\sum_{i=1}^{n}\quantity(x_{i}-\overline{x})^{T}\quantity(y_{i}-\overline{y})$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(x_{i}^{T}y_{i}-x_{i}^{T}y_{j})=\frac{1}{n^{2}}\sum_{u=1}^{m}\sum_{v=1}^{m}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{v}}(x_{i}^{T}y_{i}-x_{i}^{T}y_{j}),$ (167)
    $=e_{x}^{T}\quantity(\frac{1}{n^{2}}\sum_{u=1}^{m}n_{u}(n_{u}-1)(G_{A}^{u}-G_{B}^{u})+\frac{1}{n^{2}}\sum_{u=1}^{m}\sum_{v\neq u}n_{u}n_{v}(G_{A}^{u}-G_{E}^{uv}))e_{y},$ (168)
@@
    $G_{D,\mathcal{T}}=\frac{1}{n^{2}}\quantity(\sum_{u=1}^{m}n_{u}(n_{u}-1)(G_{A}^{u}-G_{B}^{u})+\sum_{u=1}^{m}\sum_{v\neq u}n_{u}n_{v}(G_{A}^{u}-G_{E}^{uv}))$ (170)
@@
Let $\mathcal{F}_{u}$ be a set of functions for which there exist linearly Gram-representable interpolation constraints. When considering symmetric agent-class PEP solutions $(F^{s}_{\mathcal{T}}-G^{s}_{\mathcal{T}})$ (105)-(110), the set of interpolation constraints for $\mathcal{F}_{u}$, applied independently to each agent of a given equivalence class $\mathcal{V}_{u}$, can be expressed with $f_{A}^{u}$ and $G_{A}^{u}$.
@@
The constraint $f_{i}\in\mathcal{F}_{u}$ for each $i\in\mathcal{V}_{u}$ can be written with a set of Gram-representable interpolation constraints involving function values of agent $i$ and scalar products between quantities (i.e. points and gradients) related to agent $i$. These are thus single-agent constraints, that are identical for all $i\in\mathcal{V}_{u}$ when we restrict to symmetrized agent-class solution $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (105)-(110). Therefore, they can all be expressed with $f_{A}^{u}$ and $G_{A}^{u}$ using (158) and (161). ∎
@@
Concerning the consensus steps, Theorem 9 applies to the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, defined in Section 2. Necessary interpolation constraints for this set of matrices are given in (50), (51) and (52), from Theorem 4 applied to the consensus subspace $\mathcal{C}$. Proposition 13 below shows that these interpolation constraints can be expressed in terms of $G_{C,\mathcal{T}}$ (166) and $G_{D,\mathcal{T}}$ (170), and so in terms of $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,m$).
@@
The averaging matrix interpolation constraints (50), (51) and (52), from Theorem 4 applied to the consensus subspace $\mathcal{C}$ can be expressed with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,m$), as:
@@
Let $x^{*}$ be an optimal point for the decentralized optimization problem (1). The definition constraints for the common variable $x^{*}$ in a PEP restricted to symmetrized agent-class solutions $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (105)-(110), given by the system-level stationarity constraint
@@
can be expressed with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,m$) as
@@
We consider a PEP for distributed optimization, restricted to symmetrized agent-class solutions $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (105)-(110). The definition constraints for $\overline{x}$
@@
can be written with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,m$) as
@@
We consider a PEP for distributed optimization, restricted to symmetrized agent-class solutions $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (105)-(110). Let $\mathcal{V}_{1}=\{1\}$ be one of the $m$ equivalence classes of the PEP. If $x_{1}$ is a common variable used by all the agents in the PEP, its definition constraints are given by
@@
and can be written with $G_{A}^{u}$ and $G_{E}^{uv}$ ($u,v=1,\dots,m$) as
$e_{x_{c}}^{T}(G_{A}^{u})e_{x_{c}}+e_{x}^{T}(G_{A}^{1})e_{x}-2e_{x_{c}}^{T}(G_{E}^{u1})e_{x}=0,\qquad\text{for all $u=2,\dots,m$,}$ (178)
@@
This constraint can be expressed as (178), by using relation (161) for the two first terms and definition of $G_{E}^{uv}$ (110) for the third term. ∎
@@
Let $x^{*}\in\mathbb{R}^{d}$ denote the optimal solution of the distributed optimization problem, $x_{i}^{0}\in\mathbb{R}^{d}$ the initial iterate of agent $i$, $\overline{x}^{0}$ the average initial iterate and $R\in\mathbb{R}$ a constant. When the PEP is restricted to symmetric agent-class solutions $(F^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (105)-(110), here is a list of initial conditions that can be expressed with $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ (for $u,v=1,\dots,m$):

### v2

@@ 7 Conclusion
  • [6] Sébastien Colla. Computer-Aided Analysis of Decentralized Optimization Methods. PhD thesis, 2024.
@@
  • [27] Wei Shi, Qing Ling, Kun Yuan, Gang Wu, and Wotao Yin. On the linear convergence of the ADMM in decentralized consensus optimization. IEEE Transactions on Signal Processing, 62, 07 2013.
@@
  • [30] A. B. Taylor, J. M. Hendrickx, and F. Glineur. Exact worst-case performance of first-order methods for composite convex optimization. SIAM Journal on Optimization, 27, 2017.
@@
  • [32] A. B. Taylor, J. M. Hendrickx, and F. Glineur. Smooth strongly convex interpolation and exact worst-case performance of first-order methods. Mathematical Programming, 161, 01 2017.
@@ A.1 When all agents are equivalent
Theorem 7 characterizes the settings in which we can formulate the performance computation of a distributed optimization algorithm as an SDP PEP which is totally independent of the number of agents $n$. In particular, we need all the agents to be equivalent (Assumption 1), so we can work with a fully symmetric solution $(\mathbf{f}^{s},G^{s})$ (Corollary 6.1), composed of small repeating blocks $f_{A}$, $G_{A}$, and $G_{B}$. The resulting agent-dependent PEP, expressed with blocks $f_{A}$, $G_{A}$, and $G_{B}$ has a size independent of the number of agents $n$, but the value of $n$ can still appear in the PEP and impact its solution, as any other parameter of the problem. Using appropriate changes of variables, we can obtain a PEP formulation that is fully independent of $n$, under the condition stated in part 2 of Theorem 7. In this appendix, we show how we can explicitly build this agent-independent PEP formulation for many usual settings. For that purpose, we pass through the main components of a PEP, that we have introduced in Section 2, and detail how they can be written with the blocks $f_{A}$, $G_{A}$, and $G_{B}$ and which changes of variable get rid of $n$.
@@
Let $\mathcal{F}$ be a set of functions for which there exist linearly Gram-representable interpolation constraints. In a PEP for a distributed optimization method, with fully symmetrized solutions $(\mathbf{f}^{s},G^{s})$ (76)-(77), the constraints $f_{i}\in\mathcal{F}$, for all $i\in\mathcal{V}$ can be expressed using a set of interpolation constraints that are agent-independently Gram-representable and that only involves $f_{A}$ and $G_{A}$.
@@
Each constraint $f_{i}\in\mathcal{F}$ can be written with a set of Gram-representable interpolation constraints involving function values of agent $i$ and scalar products between quantities (i.e. points and gradients) related to agent $i$. These are thus single-agent constraints that are all identical when we restrict to agent-symmetric PEP solution $(\mathbf{f}^{s},G^{s})$ (76)-(77). Therefore, they can all be expressed with $f_{A}$ and $G_{A}$ using (91) and (92). ∎
@@
Concerning the consensus steps, Theorem 7 applies to the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, defined in Section 2. Necessary interpolation constraints for this set of matrices are given in (65), (66) and (67), from Corollary 4.1. Proposition 13 below shows that these interpolation constraints are agent-independently Gram-representable, by expressing them in terms of $G_{C}$ and $G_{D}$. In principle, the compact PEP formulation could apply to a general set of averaging matrices $\mathcal{M}$ for which there exist Gram-representable interpolation constraints. Future work may include the description of other classes of averaging matrices, in particular, the extension to non-symmetric stochastic matrices with a bound on their singular value.
@@
The averaging matrix interpolation constraints (65), (66) and (67), from Corollary 4.1, are agent-independently Gram-representable and can be expressed using $G_{C}$ and $G_{D}\in\mathbb{R}^{p\times p}$ as :
@@
Firstly, we express (65) as (129). Equation (65) $\overline{X}=\overline{Y}$ means that the agent average is preserved and can be written with scalar products as :
@@
Secondly, we express (49) as (130). Equation (49) involve centered matrices $X_{\perp},Y_{\perp}\in\mathbb{R}^{nd\times t}$,
@@
$\frac{1}{n}\sum_{i=1}^{n}(U_{i}-\overline{U})^{T}(V_{i}-\overline{V})=e_{r}^{T}G_{D}e_{V}=(e_{Y}-\lambda^{-}e_{X})^{T}G_{D}(e_{Y}-\lambda^{+}e_{X})\preceq 0$
@@
Points common to all agents ($^{*}, x¯\overline{x}$, etc)
@@
[…]
e.g. $P_{i}=\quantity[y_{i}^{k}\penalty 10000\ x_{i}^{k}\penalty 10000\ g_{i}^{k}]_{k\in I}$, and the Gram matrix $G\in\mathbb{R}^{np\times np}$ is thus defined as
@@
[…]
@@
Let $x^{*}$ be an optimal point for the decentralized optimization problem (1). The definition constraints for the common variable $x^{*}$ in a PEP restricted to fully symmetric solutions $(\mathbf{f}^{s},G^{s})$ (76)-(77), given by the system-level stationarity constraint
@@
We consider a PEP for distributed optimization, restricted to fully symmetric solutions $(\mathbf{f}^{s},G^{s})$ (76)-(77). The definition constraints for $\overline{x}$
@@
$\frac{1}{n}\sum_{i=1}^{n}\quantity(f_{i}(\overline{x}^{0})-f_{i}(x^{*}))$ $\leq R$ $\hskip-113.81102pt\quantity(e_{f(\overline{x}^{0})}-e_{f(x^{*})})^{T}f_{A}\quantity(e_{f(\overline{x}^{0})}-e_{f(x^{*})})$ $\leq R$ (152)
@@
To obtain the first two reformulations, we can use relation (89) applied respectively to vectors $x_{i}^{0}-x^{*}$ and $\nabla f_{i}(x_{i}^{0})$
@@ A.2 With multiple equivalence classes of agents
Let $\mathcal{T}$ be a partition of the agent set $\mathcal{V}$ into $r$ equivalence classes of agents (see Definition 9). Considering the PEP restricted to symmetrized agent-class solutions $G^{s}_{\mathcal{T}}$ and $\mathbf{f}^{s}_{\mathcal{T}}$ (106)-(107), we now detail how we can express all the PEP elements in terms of $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,r$), to be able to use this blocks as variable of the compact SDP PEP formulation. For the SDP condition $G^{s}\succeq 0$, we can be express it with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ using Lemma 9.1. For the other PEP elements, we need to adapt the reformulation techniques (88), (89), (90), (91), (92) and (126) to the situation where the agent set $\mathcal{V}$ is partitioned into $r$ subsets of equivalent agents $\mathcal{T}=\{\mathcal{V}_{1},\dots,\mathcal{V}_{r}\}$. The terms that we can be expressed in terms of $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ are the following (for any variables $x_{i}$ and $y_{j}$ related to agents $i$ and $j$, including the case $x_{i}=y_{j}$) :
@@
    where the first equality holds because $f_{i}(x_{i})=f_{j}(x_{j})$ for any $i,j\in\mathcal{V}_{u}$, by definition of the symmetric agent-class solution $\mathbf{f}^{s}_{\mathcal{T}}$ (106). This implies that
    $\frac{1}{n}\sum_{i=1}^{n}f_{i}(x_{i})=\frac{1}{n}\sum_{u=1}^{r}\sum_{j\in V_{u}}^{n}f_{j}(x_{j})=\frac{1}{n}\sum_{u=1}^{r}n_{u}e_{f(x)}^{T}f_{A}^{u}=e_{f(x)}^{T}f_{A,\mathcal{T}},$ (156)
@@
    $\frac{1}{n}\sum_{i=1}^{n}x_{i}^{T}y_{i}=\frac{1}{n}\sum_{u=1}^{r}\sum_{j\in V_{u}}^{n}x_{j}^{T}y_{j}=\frac{1}{n}\sum_{u=1}^{r}n_{u}e_{x}^{T}G_{A}^{u}e_{y}=e_{x}^{T}(G_{A,\mathcal{T}})e_{y},$ (159)
@@
    $G_{A,\mathcal{T}}=\frac{1}{n}\sum_{u=1}^{r}n_{u}G_{A}^{u}.$ (160)
  • The average over the $n^{2}$ pairs of agents of the scalar products between two variables, each related to any agent :
    $\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}x_{i}^{T}y_{j}$ $=\frac{1}{n^{2}}\sum_{u=1}^{r}\sum_{v=1}^{r}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{v}}x_{i}^{T}y_{j}=\frac{1}{n^{2}}\sum_{u=1}^{r}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{u}}x_{i}^{T}y_{j}+\frac{1}{n^{2}}\sum_{u=1}^{r}\sum_{v\neq u}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{v}}x_{i}^{T}y_{j},$ (161)
    $\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}x_{i}^{T}y_{j}$ $=e_{x}^{T}\quantity(\frac{1}{n^{2}}\sum_{u=1}^{r}\bigl(n_{u}G_{A}^{u}+n_{u}(n_{u}-1)G_{B}^{u}\bigr)+\frac{1}{n^{2}}\sum_{u=1}^{r}\sum_{v\neq u}n_{u}n_{v}\,G_{E}^{uv})e_{y}=e_{x}^{T}(G_{C,\mathcal{T}})e_{y},$ (162)
@@
    $G_{C,\mathcal{T}}=\frac{1}{n^{2}}\quantity(\sum_{u=1}^{r}n_{u}(G_{A}^{u}+(n_{u}-1)G_{B}^{u})+\sum_{u=1}^{r}\sum_{v\neq u}n_{u}n_{v}\,G_{E}^{uv}).$ (163)
@@
    $\frac{1}{n}\sum_{i=1}^{n}\quantity(x_{i}-\overline{x})^{T}\quantity(y_{i}-\overline{y})$ $=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}(x_{i}^{T}y_{i}-x_{i}^{T}y_{j})=\frac{1}{n^{2}}\sum_{u=1}^{r}\sum_{v=1}^{r}\sum_{i\in\mathcal{V}_{u}}\sum_{j\in\mathcal{V}_{v}}(x_{i}^{T}y_{i}-x_{i}^{T}y_{j}),$ (164)
    $=e_{x}^{T}\quantity(\frac{1}{n^{2}}\sum_{u=1}^{r}n_{u}(n_{u}-1)(G_{A}^{u}-G_{B}^{u})+\frac{1}{n^{2}}\sum_{u=1}^{r}\sum_{v\neq u}n_{u}n_{v}(G_{A}^{u}-G_{E}^{uv}))e_{y},$ (165)
@@
    $G_{D,\mathcal{T}}=\frac{1}{n^{2}}\quantity(\sum_{u=1}^{r}n_{u}(n_{u}-1)(G_{A}^{u}-G_{B}^{u})+\sum_{u=1}^{r}\sum_{v\neq u}n_{u}n_{v}(G_{A}^{u}-G_{E}^{uv}))$ (167)
@@
Let $\mathcal{F}_{u}$ be a set of functions for which there exist linearly Gram-representable interpolation constraints. When considering symmetric agent-class PEP solutions $(\mathbf{f}^{s}_{\mathcal{T}}-G^{s}_{\mathcal{T}})$ (106)-(107), the set of interpolation constraints for $\mathcal{F}_{u}$, applied independently to each agent of a given equivalence class $\mathcal{V}_{u}$, can be expressed with $f_{A}^{u}$ and $G_{A}^{u}$.
@@
The constraint $f_{i}\in\mathcal{F}_{u}$ for each $i\in\mathcal{V}_{u}$ can be written with a set of Gram-representable interpolation constraints involving function values of agent $i$ and scalar products between quantities (i.e. points and gradients) related to agent $i$. These are thus single-agent constraints, that are identical for all $i\in\mathcal{V}_{u}$ when we restrict to a symmetrized agent-class solution $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (106)-(107). Therefore, they can all be expressed with $f_{A}^{u}$ and $G_{A}^{u}$ using (155) and (158). ∎
@@
Concerning the consensus steps, Theorem 9 applies to the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$, defined in Section 2. Necessary interpolation constraints for this set of matrices are given in (65), (66) and (67), from Corollary 4.1. Proposition 13 below shows that these interpolation constraints can be expressed in terms of $G_{C,\mathcal{T}}$ (163) and $G_{D,\mathcal{T}}$ (167), and so in terms of $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,r$).
@@
The averaging matrix interpolation constraints (65), (66) and (67), from Corollary 4.1 can be expressed with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,r$), as :
@@
Let $x^{*}$ be an optimal point for the decentralized optimization problem (1). The definition constraints for the common variable $x^{*}$ in a PEP restricted to symmetrized agent-class solutions $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (106)-(107), given by the system-level stationarity constraint
@@
can be expressed with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,r$) as
@@
We consider a PEP for distributed optimization, restricted to symmetrized agent-class solutions $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (106)-(107). The definition constraints for $\overline{x}$
@@
can be written with $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,r$) as
@@
We consider a PEP for distributed optimization, restricted to symmetrized agent-class solutions $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (106)-(107). Let $\mathcal{V}_{1}=\{1\}$ be one of the $r$ equivalence classes of the PEP. If $x_{1}$ is a common variable used by all the agents in the PEP, its definition constraints are given by
@@
and can be written with $G_{A}^{u}$ and $G_{E}^{uv}$ ($u,v=1,\dots,r$) as
$e_{x_{c}}^{T}(G_{A}^{u})e_{x_{c}}+e_{x}^{T}(G_{A}^{1})e_{x}-2e_{x_{c}}^{T}(G_{E}^{u1})e_{x}=0,\qquad\text{for all $u=2,\dots,r$,}$ (175)
@@
This constraint can be expressed as (175), by using relation (158) for the first two terms and definition of $G_{E}^{uv}$ (107) for the third term. ∎
@@
Let $x^{*}\in\mathbb{R}^{d}$ denote the optimal solution of the distributed optimization problem, $x_{i}^{0}\in\mathbb{R}^{d}$ the initial iterate of agent $i$, $\overline{x}^{0}$ the average initial iterate and $R\in\mathbb{R}$ a constant. When the PEP is restricted to symmetric agent-class solutions $(\mathbf{f}^{s}_{\mathcal{T}},G^{s}_{\mathcal{T}})$ (106)-(107), here is a list of initial conditions that can be expressed with $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ (for $u,v=1,\dots,r$) :